OpenStudy (anonymous):

Plzz help! Choose the correct description of the graph of the compound inequality x - 2 > -4 and 3x less than or greater to 15

OpenStudy (anonymous):

A. A number line with an open circle on -2, shading to the left, and a closed circle on 5, shading to the right B. A number line with a closed circle on -2, shading to the left, and an open circle on 5, shading to the right C. A number line with an open circle on -2, a closed circle on 5, and shading in between D. A number line with a closed circle on -2, an open circle on 5, and shading in between

OpenStudy (anonymous):

@sparklingbubbles can u help?

OpenStudy (anonymous):

@iGreen @BloomLocke367

OpenStudy (anonymous):

can u help? ^^

OpenStudy (igreen):

Simplify it first

OpenStudy (igreen):

x - 2 > -4 Add 2 to both sides

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it would be 2

OpenStudy (anonymous):

2-* hah sorry

OpenStudy (anonymous):

-2

OpenStudy (igreen):

Yep. x > -2 \(3x \le 15\) Divide 3 to both sides

OpenStudy (anonymous):

thx i have to go

OpenStudy (igreen):

Np

OpenStudy (anonymous):

@inowalst @Nnesha @matt101

OpenStudy (matt101):

What do you need help with?

OpenStudy (inowalst):

^

OpenStudy (anonymous):

@iGreen @inowalst can you guys help me finish it plz hah ^

OpenStudy (igreen):

Okay, \(3 \le 15\) Divide 3 to both sides, 15 / 3 = ?

OpenStudy (igreen):

That's 3x..sorry.

OpenStudy (anonymous):

= 5

OpenStudy (anonymous):

it's cool

OpenStudy (igreen):

Yep, so we have: \(x \le 5\) and \(x > -2\)

OpenStudy (igreen):

Do you know what those would look like on a graph?

OpenStudy (anonymous):

well seance it's a negative 2 do we switch the sighs rights

OpenStudy (anonymous):

> to this < ?

OpenStudy (igreen):

No..

OpenStudy (igreen):

That's only when we divide or multiply by a negative number..but it's already simplified

OpenStudy (anonymous):

ohh ok so would the answer be C?

OpenStudy (igreen):

Yes!

OpenStudy (anonymous):

thank you :)